OpenStudy (anonymous):

Three pizzerias claim they offer the pizza with the largest circumference. Each of the 8 slices of Mario's Pizzeria's large pizza has an area of 14.13 square inches. Each of the 16 slices of Tito's Pizzeria's large pizza has an area of 12.56 square inches. Each of the 10 slices of Peter's Pizzeria's large pizza has an area of 15.386 square inches. Place the pizzas in order of least to greatest circumference. Assume all pizzas are circles and assume all pizzas have the same thickness. Use 3.14 for .

OpenStudy (mertsj):

Area of first one: 8(14.13)=113.04 \[A=\pi r^2\] \[113.04=\pi r^2\] \[\frac{113.04}{\pi}=r^2=36\] r=6

OpenStudy (mertsj):

Use that idea to find the radius of each pizza. The pizza with the smallest radius has the smallest circumference.

OpenStudy (anonymous):

ok ty

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

wait why does it say math prossesing error?

OpenStudy (anonymous):

@Mertsj

OpenStudy (mertsj):

What says math processing error?

OpenStudy (anonymous):

under the area of first one everything else says math prossesing erroe

OpenStudy (anonymous):

error*

OpenStudy (mertsj):

Maybe the equation editor is not working.

OpenStudy (mertsj):

The area of the first pizza is 113.04 Formula for area of circle: A=pir^2 So 113.04=3.14 r^2 Divide both sides by 3.14 and you see that 36 = r^2 and so radius = 6

OpenStudy (mertsj):

Do you understand?

OpenStudy (anonymous):

o ok thank u so much

OpenStudy (mertsj):

yw